I'm not sure what to do with the first part of this problem can I get the answer so I can work backwards and figure it out?
1-(-9n-1)=7(-n+2)+1
n=?

Answer by unlockmath(1688)

Hello,
1-(-9n-1)=7(-n+2)+1
Rewritten as:
1+9n+1=-7n+14+1
combine like terms:
9n+2=-7n+15
Add 7n and subtract 2 to get:
16n=13
Divide by 16
n=13/16
Make sense?
RJ
www.maht-unlock.com